Mason Company has two manufacturing departments—Machining and Assembly. The company considers all of its manufacturing overhead costs to be fixed costs. It provided the following estimates at the beginning of the year as well as the following information with respect to Jobs A and B:

Estimated Data Machining Assembly Total
Manufacturing overhead $ 20,184,000 $ 841,000 $ 21,025,000
Direct labor hours 29,000 696,000 725,000
Machine hours 696,000 24,000 720,000
Job A Machining Assembly Total
Direct labor hours 5 10 15
Machine hours 11 2 13
Job B Machining Assembly Total
Direct labor hours 4 5 9
Machine hours 12 3 15

Required:

1. If Mason Company uses a plantwide predetermined overhead rate with direct labor-hours as the allocation base, how much manufacturing overhead cost would be applied to Job A? Job B? (Round your answers to the nearest whole dollar amount.)

2. Assume that Mason Company uses departmental predetermined overhead rates. The Machining Department is allocated based on machine-hours and the Assembly Department is allocated based on direct labor-hours. How much manufacturing overhead cost would be applied to Job A? Job B? (Round your intermediate calculations and final answers to 2 decimal places.)
Answer #1
Required 1 :
Plantwide predetermined overhead rate = Total estimated manufacturing overhead / Total estimated direct labor hours = 21025000 / 725000 29
Job A Job B

Manufacturing overhead cost applied

[ Direct labor hours * Plantwide predetermined overhead rate ]

435

[ 29*15 ]

261

[ 29*9 ]

Required 2 :
Machining department overhead rate = Estimated manufacturing overhead of machining department / Estimated machine hours in machining department = 20184000 / 696000 29
Assembly department overhead rate = Estimated manufacturing overhead in Assembly department / Estimated direct labor hours in assembly department = 841000 / 696000 1.21
Job A Job B

Overhead related to Machining department

[ Machine hours in machining department * Machining department overhead rate ]

319

[ 29*11 ]

348

[ 29*12 ]

Overhead related to assembly department

[ Direct labor hours in Assembly department * Assembly department overhead rate ]

12.10

[ 1.21*10 ]

6.05

[ 1.21*5 ]

Manufacturing overhead cost applied 331 354